SpaceRef

SpaceRef


ESA Venus Express Status Report No. 53 - End of Conjuntion Operations

Status Report From: ESA Venus Express Mission
Posted: Thursday, November 16, 2006

image

13 Nov 2006 10:14

Report for Period 5 November to 11 November 2006

During the reporting period the mission operations have been conducted according to the plan. The solar conjunction ended on 9 November. All data collected during the conjunction period were dumped, the spacecraft re-configured for routine mission and science operations were resumed in orbit #204.

The table below shows a chronology of the main activities in the reporting period:

MET (Day) Date DOY VPER# Main Activity
362 05/10/06 309 198 Conjunction operations
363 06/10/06 310 199 Conjunction operations
364 07/10/06 311 200 Conjunction operations
365 08/11/06 312 201 Conjunction operations
Increasing TM/TC bitrates
366 09/11/06 313 202 Configure S/C for routine operations,
Dump S/C telemetry from conjunction period (1st part),
Upload schedule to resume science operations
367 10/11/06 314 203 Dump S/C telemetry from conjunction period (2nd part)
368 11/11/06 315 204 Routine Mission operations

At the end of the last Cebreros pass in the reporting period (DOY 315, 18:00z) Venus Express was orbiting Venus at 256 million km from the Earth. The one-way signal travel time was 852 seconds.


Science Update

The following set of images has been released since the previous status report.

13 November 2006
VIRTIS - Visible and Infrared Thermal Imaging Spectrometer

Date: 23 September 2006

Description: False-colour image of area close to the south pole imaged at 1.7 μm, revealing the differences in cloud cover thickness and the thermal emission coming through the clouds.

Date: 29 July 2006

Description: Monochrome image of the south polar region imaged at 1.7 μm, showing the distinctive pattern of cloud wave structure at around 55° southern latitude.

 


Payload Activities

ASPERA
Routine operation of the instrument was resumed on DOY 314.

MAG
Routine operation of the instrument was resumed on DOY 314.

PFS
The instrument is OFF and was not operated.

SPICAV
Routine operation of the instrument was resumed on DOY 314.

VeRA
The USO is nominally powered but muted. No RSI experiments were performed during the reporting period.

VIRTIS
Routine operation of the instrument was resumed on DOY 314.
A workaround to avoid HK bursts, which recently led to spacecraft safe modes, was tested and approved by the PI. After this the on-board execution of VIRTIS TCs was enabled again and the instrument was operated successfully in orbit 204.

VMC
Routine operation of the instrument was resumed on DOY 314.

 

Mission Planning

Short Term Planning
Operations of STP027 were executed successfully and the planning of STP028 was completed in time.

The pre-check of PORs for STP029 has been finished with success.

Medium Term Planning
MTP010 planning is on-going.

Long Term Planning
Discussions on the data amount produced by the movie observation, to be performed in April 2007, are ongoing.


Future Milestones

Swap from HGA1 to HGA2 in May 2007. An operational strategy to avoid VMC illumination is under discussion and has to be agreed with Astrium and Project.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks